OpenStudy (anonymous):

2x^2+3=2x 2x^2+3=2x @Mathematics

OpenStudy (anonymous):

Solve completely with Complex numbers

OpenStudy (mathmagician):

\[2x^2-2x+3=0\] \[D=4-4*3*2=-20\] \[x1,2=(2+-\sqrt{-20})/4=(1+-\sqrt{5i^2})/2=(1+-i \sqrt{5})/2\]

OpenStudy (anonymous):

So the answer isn't: \[x=\frac{1}{2} \pm 1 \frac{\sqrt{5}}{2}i\]

OpenStudy (mathmagician):

its the same as i wrote, just written in other form

OpenStudy (anonymous):

Ahh... OK couldn't tell, I was confused.